\begin{thebibliography}{99}
%\bibitem{AP} D.~Abramovich, A.~Polishchuk,
%{\it Sheaves of t-structures and valuative
%criteria for stable complexes}, J. Reine Angew. Math.  590 (2006), 89--130.
%\bibitem{Atiyah} Atiyah, {\it Complex analytic connections in fibre bundles}
\bibitem{AC} R.~Auffarth, G.~Codogni, {\it Theta divisors whose Gauss map has a fiber of positive dimension}, J. Algebra 548 (2020), 153--161.
\bibitem{BGP} V.~Baranovsky, V.~Ginzburg, J.~Pecharich, {\it Deformations of line bundles on coisotropic subvarieties},
Q. J. Math. 63 (2012), no. 3, 525--537.
\bibitem{BGKP} V.~Baranovsky, V.~Ginzburg, D.~Kaledin, J.~Pecharich,
{\it Quantization of line bundles on lagrangian subvarieties},
Selecta Math. (N.S.) 22 (2016), no. 1, 1--25.
\bibitem{BP} A.~Beilinson, A.~Polishchuk,
{\it Torelli theorem via Fourier-Mukai    transform}, in {\it Moduli of abelian varieties}, C.~Faber et al. ed.,
         127--132, Birkh\"auser, 2001. 
\bibitem{BBP} O.~Ben-Bassat, J.~Bloch, T.~Pantev, {\it Non-commutative tori and Fourier-Mukai duality},
Compos. Math. 143 (2007), no. 2, 423--475.
%\bibitem{Cortinas} Cortinas
\bibitem{DP} B.~Dyer, A.~Polishchuk,
{\it NC-smooth algebroid thickenings for families of vector bundles and quiver representations}, 
     Compositio Math. 155 (2019), 681--710.
\bibitem{Fedosov} B.~Fedosov,
{\it A simple geometrical construction of deformation quantization},
J. Differential Geom. 40 (1994), no. 2, 213--238.
%\bibitem{CQ} Cuntz, Quillen
\bibitem{Kapranov} M.~Kapranov, {\it Noncommutative geometry based on commutator expansions},
J. Reine Angew. Math. 505 (1998), 73--118.
%\bibitem{Kapranov2} M.~Kapranov, {\it Rozansky-Witten invariants via Atiyah classes}
%\bibitem{KM} I.~Kriz, J.~P.~May, {\it Operads, algebras, modules, and motives}.
%\bibitem{Laumon} Laumon
\bibitem{MMG} A.~C.~L\'opez Mart\'in, E.~Mistretta, D. S\'anchez G\'omez,
{\it A Characterization of Jacobians by the Existence of Picard Bundles}, Le Mathematique, Vol. LXIII (2008), Fasc. I, 167--179.
\bibitem{Mukai} S.~Mukai,  {\it Duality between $D(X)$ and $D(\hat{X})$
with its application to Picard sheaves}, Nagoya Math. J. 81 (1981), 153--175.
\bibitem{Mukai-2} S.~Mukai, {\it Fourier functor and its application to
the moduli of bundles on an abelian variety}, Adv. Studies in Pure Math. 10
(1987), 515--550.
\bibitem{P-Tu} A.~Polishchuk, J.~Tu,
   {\it DG-resolutions of NC-smooth thickenings and NC-Fourier-Mukai transforms}, 
   Math. Annalen 360 (2014), 79--156.
%\bibitem{Serre} Serre, {\it Galois cohomology}
\bibitem{Toda} Y.~Toda, {\it Deformations and Fourier-Mukai transforms}, J. Differential Geom. 81 (2009), no. 1, 197--224.
\end{thebibliography}

\documentclass[12pt,reqno]{amsart}
\usepackage{diagrams}
\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage[all,arc]{xy}
\usepackage{cite}
\usepackage{tikz-cd}
\usepackage{MnSymbol}

\newcommand{\Per}{\operatorname{Per}}

\newcommand{\pa}{\partial}

\newcommand{\WW}{{\mathcal W}}

\newcommand{\ssl}{{\frak sl}}

\newcommand{\ch}{\operatorname{ch}}
\newcommand{\CH}{\operatorname{CH}}
\renewcommand{\mod}{\operatorname{mod}}

\newcommand{\Cone}{\operatorname{Cone}}

\newcommand{\und}{\underline}

\newcommand{\OO}{{\mathcal O}}

\newcommand{\coker}{\operatorname{coker}}
\newcommand{\DD}{{\mathcal D}}
\newcommand{\NN}{{\mathcal N}}

\newcommand{\hra}{\hookrightarrow}
\newcommand{\lan}{\langle}
\newcommand{\ran}{\rangle}
\newcommand{\Coh}{\operatorname{Coh}}
\newcommand{\GG}{{\mathcal G}}

\newcommand{\Spec}{\operatorname{Spec}}
\newcommand{\Sing}{\operatorname{Sing}}
\newcommand{\Td}{\operatorname{Td}}

\newcommand{\Th}{\Theta}
\newcommand{\si}{\sigma}

\newcommand{\Pic}{\operatorname{Pic}}

\newcommand{\de}{\delta}

\renewcommand{\ker}{\operatorname{ker}}

\numberwithin{equation}{subsection}

\newtheorem{thm}{Theorem}[subsection]
\newtheorem{prop}[thm]{Proposition}

\newtheorem{lem}[thm]{Lemma}
\newtheorem{cor}[thm]{Corollary}
{  \theoremstyle{definition}

\newtheorem{exs}[thm]{Examples}
\newtheorem{rem}[thm]{Remark}
\newtheorem{rems}[thm]{Remarks}
}

\newcommand{\Pf}{\noindent {\it Proof}}
\newcommand{\id}{\operatorname{id}}

\newcommand{\LLie}{{\cal L}ie}
\newcommand{\ov}{\overline}
\newcommand{\we}{\wedge}

\renewcommand{\AA}{{\mathcal A}}

\newcommand{\FF}{{\mathcal F}}

\newcommand{\JJ}{{\mathcal J}}

\newcommand{\Om}{\Omega}

\newcommand{\Hom}{\operatorname{Hom}}

\newcommand{\Ext}{\operatorname{Ext}}

\renewcommand{\a}{\alpha}
\renewcommand{\b}{\beta}
\newcommand{\om}{\omega}

\newcommand{\la}{\lambda}
\renewcommand{\th}{\theta}
\newcommand{\C}{{\Bbb C}}

\newcommand{\Z}{{\Bbb Z}}
\newcommand{\Q}{{\Bbb Q}}

\newcommand{\wt}{\widetilde}
\newcommand{\ot}{\otimes}

\newcommand{\sub}{\subset}
\newcommand{\ed}{\qed\vspace{3mm}}

\newcommand{\Qcoh}{\operatorname{Qcoh}}

\newcommand{\NS}{\operatorname{NS}}

\newcommand{\cl}{\operatorname{cl}}

\title{NC-smooth thickenings and Jacobians}
\author{Alexander Polishchuk}
\address{
    Department of Mathematics, 
    University of Oregon, 
    Eugene, OR 97403, USA; and National Research University Higher School of Economics, Moscow, Russia
  }
 \email{apolish@uoregon.edu}

\thanks{Supported in part by NSF grant DMS-2349388, by the Simons Travel grant MPS-TSM-00002745,
and within the framework of the HSE University Basic Research Program.}

\begin{document}

\begin{abstract}
We prove that a bounded complex of coherent sheaves on an abelian variety $A$, whose Fourier-Mukai transform has support of dimension $\le 1$, 
extends to a perfect complex on the standard NC-smooth thickening of $A$ and on the deformation quantization of any constant Poisson bracket on $A$. 
We discuss a related conjectural characterization of Jacobians in
terms of NC-smooth thickenings.
\end{abstract}

\maketitle

\centerline{\sc Introduction}

NC-smooth thickenings of smooth algebraic varieties were introduced by Kapranov \cite{Kapranov} as a way to study noncommutative geometry
in the ``formal neighborhood" of a commutative variety. The model example of an NC-smooth thickening is a completion of a free algebra in $n$ generators
with respect to the commutator filtration (see Sec.\ \ref{NC-thick-sec} below). 
Kapranov's idea is to consider sheaves of algebras $\OO_X^{NC}$ over smooth varieties that look like this model example in a formal neighborhood of each point (such that by taking the quotient of $\OO_X^{NC}$ by the ideal generated by commutators one recovers the structure sheaf $\OO_X$).
Kapranov proved existence and uniqueness of such NC-smooth thickenings for smooth affine varieties, and constructed some non-affine examples.
Non-affine examples were further studied in \cite{P-Tu} and \cite{DP}. 

Every abelian variety $A$ possesses a natural translation-invariant NC-smooth thickening $A^{NC}$ that we will call {\it standard}. 
In holomorphic terms, this can be obtained by the descent of the standard
NC-smooth thickening of the affine space with respect to translations by a lattice. Over any field of characteristic zero, there exists an algebraic construction using
a canonical flat connection on $A$ (see \cite{P-Tu}). In the case when $A=J$, the Jacobian of a smooth projective curve $C$, there is a way to think of the standard NC-smooth thickening of $J$ as representing a natural noncommutative moduli functor (see \cite[Sec.\ 4]{P-Tu}), and this gives rise to an NC-smooth analog of the Fourier-Mukai transform
that produces sheaves on $J^{NC}$ from sheaves on $C$ (but it is far from being an equivalence).

In this paper we show that the theta divisor $\Th$ on $J$ (defined up to a translation) also extends in some sense to the standard NC-smooth thickening $J^{NC}$. Namely, we prove that there exists a coherent sheaf supported on $\Th$, generically of rank $1$ on $\Th$, which extends to an object of the derived category of $J^{NC}$ represented by a bounded complex of vector bundles 
(see Theorem \ref{Jac-thm}). We conjecture that this property characterizes Jacobians among principally polarized varieties with irreducible theta divisors.


\medskip

\noindent
{\bf Conjecture A}.
{\it Let $(A, \Theta)$ be a principally polarized abelian variety with its theta divisor, where
$\Theta$ is irreducible.
Then it is the Jacobian of a smooth curve if and only if there exists a coherent sheaf $\FF$ on $A$, supported on $\Theta$ and of length $1$ at the general point of $\Theta$,
such that $\FF$ extends to an object of $D^b(A^{NC})$ represented by a bounded complex of vector bundles
(where $A^{NC}$ is the standard  NC-smooth thickening of $A$).
}

\medskip

The ``only if" part is Theorem \ref{Jac-thm}. As an evidence for the ``if" part, we prove that existence of a coherent sheaf $\FF$ as in Conjecture A implies that
$(A,\Theta)$ belongs to the Andreotti-Mayer locus, i.e., that the dimension of the singular locus of $\Theta$ is $\ge g-4$, where $g=\dim A$ (see Cor.\ \ref{ch-2-cor}).
In fact, to deduce this we only need an extension to the standard 1-smooth thickening of $A$ (see Sec.\ \ref{NC-thick-sec}).

As a possible approach to Conjecture A, we suggest the following conjecture giving a criterion of extendability to the standard NC-smooth thickening in terms of
the Fourier-Mukai transform $F:D^b(A)\to D^b(\hat{A})$ (see \cite{Mukai}). 

\medskip

\noindent
{\bf Conjecture B}.
{\it Let $A$ be an abelian variety, $A^{NC}$ its standard NC-smooth thickening. An object $E\in D^b(A)$ extends to a bounded complex of vector bundles on $A^{NC}$
if and only if the Fourier-Mukai transform of $E$ is supported in dimension $\le 1$.
}

\medskip

We prove the ``if" part in Theorem \ref{NC-FM-thm}. 
Note that existence of a resolution by locally free $\OO_{A^{NC}}$-modules of finite rank is much more restrictive than in commutative case (see e.g., Cor.\ \ref{codim-cor}).

Instead of asking which sheaves on a variety $X$ extend to an NC-smooth thickening, one can ask a similar question for deformation quantizations of Poisson structures on $X$.
In the case of constant Poisson structures on abelian varieties and the corresponding Moyal deformation quantizations, we show that there is a relation between these questions.
Namely, we construct a homomorphism from the structure sheaf of standard NC-smooth thickening $\OO_A^{NC}$ to the deformation quantization algebra associated with any constant Poisson bivector (see Proposition \ref{NC-to-quant-prop}). 
Using the result of \cite{BGP} we prove that existence of an extension of an appropriate sheaf on the theta-divisor to the first-order deformation
quantization associated with sufficiently generic constant Poisson bivectors implies that the dimension of the singular locus of $\Theta$ is $\ge \dim A-4$ (see Prop.\ \ref{1st-order-def-qu-prop}).

On the other hand, we show that the ``only if" part of Conjecture B does not hold if we replace the NC-smooth thickening by the corresponding first-order thickening.
Namely, using the results of Toda \cite{Toda} we give an example of a coherent sheaf on an abelian surface which extends to the standard 1-smooth thickening but 
whose Fourier-Mukai transform has full support (see Sec.\ \ref{surface-ex-sec}). 



\bigskip

\noindent
{\it Conventions and notation}. We work over $\C$.
We denote by $\ch^H$ and $c_1^H$ the characteristic classes with values in $H^*(X,\Q)$ (sometimes we also view $c_1^H$ as an element
of $H^1(X,\Om^1_X)$); and by
$\ch$ and $c_1$ the characteristic classes with values in the Chow group. For a smooth variety $X$ we denote by $D^b(X)$ the bounded derived category of coherent
sheaves on $X$.

\section{NC-smooth thickenings and NC-Fourier transform}

\subsection{NC-smooth thickenings}\label{NC-thick-sec}

We refer to \cite{Kapranov} and \cite[Sec.\ 2.1]{P-Tu} for basics on NC-smooth thickenings.

The commutator filtration $F^dR$ on a ring $R$ is defined by letting $F^dR$ be a two-sided ideal generated by all expressions containing $d$-commutators
(possibly nested). The ring is called NC-nilpotent if $F^dR=0$ for some $d$, and NC-complete if the natural map $R\to\varprojlim R/F^dR$ is an isomorphism.
NC-smoothness for NC-complete algebras is defined similarly to formal smoothness, in terms of nilpotent extensions of NC-smooth algebras.

These notions can be globalized, so one can consider sheaves of NC-smooth algebras. For a smooth variety $X$, an NC-smooth thickening of $X$ is defined
as a sheaf of NC-smooth algebras $\OO_X^{NC}$ together with an isomorphism $\OO_X^{NC}/F^1\simeq \OO_X$. 

One can also consider $n$-smooth thickenings by considering only NC-algebras (or sheaves) with $F^{n+1}=0$. For example, a $1$-smooth thickening of a smooth variety
$X$ is a central square zero extension of sheaves of algebras 
$$0\to \NN\to \wt{\OO}_X\to \OO_X\to 0$$
such that the commutator pairing $f,g\mapsto [\wt{f},\wt{g}]$, where $\wt{f}$ and $\wt{g}$ are liftings of $f,g\in \OO_X$ to $\wt{\OO}_X$, induces an isomorphism
$\Om_X^2\rTo{\sim} \NN$. In particular, for every $X$ there is a {\it standard} 1-smooth thickening, 
$$\OO_X^{1-NC}=\Om_X^2\oplus \OO_X,$$
equipped with the product $(\a,f)\cdot (\b,g)=(g\a+f\b+dg\we dg,fg)$
(see \cite[(1.3.9)]{Kapranov}).

In \cite{P-Tu} we developed a construction of an NC-smooth thickening associated with a torsion free connection on the tangent bundle $T_X$ (in the case of a flat connection it is recalled in Sec.\ \ref{dg-constr-sec} below).
Applying this construction to the natural flat connection on an abelian variety $A$ we get the {\it standard} NC-smooth thickening $A^{NC}$
(more generally, this works for commutative algebraic groups, see \cite[Def.\ 2.3.19]{P-Tu}). The induced 1-smooth thickening $\OO_A^{NC}/F^2$ is the standard
1-smooth thickening of $A$ discussed above (by \cite[Prop.\ 2.3.18]{P-Tu}).

\subsection{NC-Fourier transform}

Let $A$ be an abelian variety, $\hat{A}$ the dual abelian variety.
It was observed in \cite{P-Tu} that the restriction of Fourier-Mukai transform $F:D^b(\hat{A})\to D^b(A)$ to $D^b(C)$, for a curve $C\sub \hat{A}$,
extends to an NC-Fourier transform $F^{NC}:D^b(C)\to \Per(A^{NC})$, where $A^{NC}$ is the standard NC-smooth thickening of $A$, and $\Per(A^{NC})$
denotes the full subcategory in the derived category of $\OO_{A^{NC}}$-modules formed by bounded complexes of vector bundles (i.e., locally free left modules of finite rank).

In the next theorem we use a slight extension of this construction to prove the ``if" part of Conjecture B.





\begin{thm}\label{NC-FM-thm} Let $A$ be an abelian variety, $A^{NC}$ its standard thickening.
If the Fourier-Mukai transform of $E\in D^b(A)$ has support of dimension $\le 1$ then
$E$ extends to a perfect object in $D^b(A^{NC})$. Furthermore, for any closed subset $Z\sub \hat{A}$ of dimension $\le 1$, 
there exists a functor 
$$D^b_Z(\hat{A})\to \Per(A^{NC}),$$
where $D^b_Z(\hat{A})\sub D^b(\hat{A})$ is the full subcategory of objects with support on $Z$, 
extending the Fourier-Mukai transform $D^b_Z(\hat{A})\to D^b(A)$.
\end{thm}

\Pf . 
Let $Z\subset \hat{A}$ be a closed subset 
of dimension $1$. We can find two affine open subsets $U_1,U_2\sub\hat{A}$
such that $Z\sub U:=U_1\cup U_2$.
We have $D^b_Z(\hat{A})=D^b_Z(U)$, so it is enough
to construct an integral transform $D^b_Z(U)\to \Per(A^{NC})$ 
extending the Fourier transform $D^b_Z(U)\to D^b(A)$.
As in \cite[Sec.\ 4.3]{P-Tu}, for this we construct an extension of the Poincare line bundle to $U\times A^{NC}$.

Let $A^\natural\to \hat{A}$ denote the $H^0(A,\Om^1)$-torsor parametrizing line bundles with connections on $A$.
Since $U_i$ are affine, we can choose liftings of $U_1$ and $U_2$ to $A^\natural$. 
As is explained in \cite[Sec.\ 4.3]{P-Tu}, using these liftings we get extensions of the Poincar\'e bundle over $U_i\times A$ to $U_i\times A^{NC}$ and
an isomorphism between the induced extensions on $(U_1\cap U_2)\times A^{NC}$.
Hence, we obtain a line bundle on $U\times A^{NC}$ that we can use as a kernel of an integral
transform 
$$F_U:D^b(\Qcoh(U))\to D(\OO^{NC}_A-\mod)$$ 
which lifts the corresponding integral transform from $D^b(\Qcoh(U)$ to $D(\Qcoh A)$.
It remains to prove that $F_U(\FF)$ is a perfect complex for any $\FF$ supported on $Z$.
It is enough to prove this for skyscraper sheaves and for coherent sheaves obtained as push-forwards
from normalizations of irreducible components of $Z$. But this follows from \cite[Thm.\ 4.3.6]{P-Tu}.
\ed



One indirect evidence if favor of Conjecture B is that both conditions imply that $\ch^H_{\ge 2}(E)=0$: 
for the condition that $F(E)$ is supported in dimension $\le 1$ this follows from the formula for the action of the Fourier-Mukai transform on cohomology
(see \cite[Cor.\ 1.18]{Mukai-2}). For the condition that $E$ extends to a perfect complex on $A^{NS}$ this is a consequence of the following result.

\begin{prop}\label{ch-prop}
Let $X$ be a smooth projective variety.
If an object $E\in D^b(X)$ extends to a bounded complex of vector bundles on the standard 1-smooth thickening of $X$ then $\ch^H_{\ge 2}(E)=0$.
\end{prop}

\begin{proof} This follows immediately from \cite[Cor.\ 3.3.5]{P-Tu}.
\end{proof}

\begin{cor}\label{codim-cor} 
Let $\FF$ be a nonzero coherent sheaf on a smooth connected projective variety $X$.

\noindent 
(i) If $\FF$ has support of codimension $\ge 2$ then $\FF$ does not extend to a perfect complex
on the standard 1-smooth thickening of $X$. 

\noindent
(ii) Assume $\FF=i_*\GG$, where $i:D\to X$ is the embedding of a smooth connected divisor, and $\GG$ is a coherent
sheaf of rank $r$ on $D$. If $\FF$ extends to a perfect complex on the standard 1-smooth thickening of $X$ then one has
$$i_*(c^H_1(\GG)-r\cdot c_1^H(\OO(D)|_D)/2)=0$$
in $H^4(X,\Q)$.
\end{cor}

\begin{proof}
(i) Let $n=\dim X$, and let $m$ be the dimension of the support of $X$. Then $\ch(\FF)_{n-m}=\sum_i a_i [Z_i]\in \CH^{n-m}(X)_{\Q}$ for some 
irreducible subvarieties $Z_i$ of dimension $m$ and some positive coefficients $a_i$, and $\ch(\FF)_{<n-m}=0$. It follows that if $h$ is an ample divisor class
then $\ch^H(\FF)_{n-m}\cdot h^m\neq 0$, so $\ch^H(\FF)_{n-m}\neq 0$. Thus, if $m\le n-2$ we cannot have $\ch^H_{\ge 2}(\FF)=0$. So, by Proposition
\ref{ch-prop}, $\FF$ cannot extend to the standard 1-smooth thickening of $X$.

\noindent
(ii) This follows immediately by applying the Grothendieck-Riemann-Roch theorem
to rewrite the condition $\ch^H_2(i_*\GG)=0$.
\end{proof}

\begin{rems} 1. Note that both conditions in Conjecture B are preserved by pull-backs under homomorphisms between abelian varieties (as well as by translations) and by tensoring with line bundles in $\Pic^0(A)$. This follows from the standard
compatibility of the Fourier-Mukai transform with these operations and from the fact that they extend to the standard NC-smooth smooth thickenings
(see \cite[Prop.\ 2.3.20]{P-Tu} and  \cite[Thm.\ 3.3.8(ii)]{P-Tu}).


\noindent
2. The results of \cite[Sec.\ 3.2]{P-Tu} suggest that a plausible analog of Conjecture B for $\OO_A^{NC}$-bimodules is that an object $E\in D^b(A)$ extends to a complex of $\OO_A^{NC}$-bimodules, 
locally free of finite rank as left or right modules, if and only the Fourier-Mukai transform of $E$ has $0$-dimensional support.
\end{rems}

\subsection{Extensions of the theta-divisor to the NC-thickening}


Using Theorem \ref{NC-FM-thm} we get a proof of the ``only if" direction of Conjecture A.

\begin{thm}\label{Jac-thm}
Let $(J,\Th)$ be the Jacobian of a smooth projective curve $C$, with its theta divisor. Then there exists a line bundle $L$ on the smooth part $\Th^{ns}$ of $\Th$,
such that the coherent sheaf $\FF:=j^{ns}_*L$ on $\Th$, where $j^{ns}:\Th^{ns}\to J$ is the natural embedding, 
extends to a perfect complex on $J^{NC}$, the standard NC-thickening of $J$.
\end{thm}

\begin{proof}
Let $i:C\to J$ be the embedding associated with some point, and let $L$ be a line bundle of degree $g-1$
on $C$. Then the Fourier-Mukai transform of $i_*L$, is of the form $\FF[-1]$, where $\FF$ is of the described form (see \cite[Prop.\ 2.3]{BP}). It remains
to apply Theorem \ref{NC-FM-thm}.
\end{proof}

\begin{rem}
Note that if a coherent sheaf $\FF$ on $X$ extends to a bounded complex $P^\bullet$ of locally free modules of finite rank over for some NC-nilpotent thickening $A\to \OO_X$
then we can assume this complex to be concentrated in non-positive degrees (i.e., $P^{>0}=0$). Indeed, if $P^{>n+1}=0$ for $n\ge 0$, then 
$\coker(d_n:P^n\to P^{n+1})$ becomes zero when tensored with $\OO_X$. Hence, $\coker(d_n)=0$, so we can replace $P^n$ with $\ker(d_n)$ and $P^{n+1}$ with $0$.
Iterating this we will get a complex in non-positive degrees.
\end{rem}

Recall (see \cite{Mukai}) that a coherent sheaf $\FF$ on an abelian variety $A$ is called an IT-sheaf (resp., WIT-sheaf) of index $i$ if $H^j(A,\FF\ot P)=0$ for $j\neq i$ for
every $P\in \Pic^0(A)$ (resp., $\und{H}^j F(\FF)=0$ for $j\neq i$, where $F:D^b(A)\to D^b(\hat{A})$ is the Fourier-Mukai transform).

The following characterization of Jacobians is due to \cite{MMG} (based on Matsusaka-Ran criterion).

\begin{thm}\label{MMG-thm} 
(\cite[Thm.\ 4.3]{MMG}) Let $(A,\Theta)$ be an indecomposable principally polarized abelian variety.
Suppose there exists a WIT-sheaf $\GG$ of index $g$ on $A$ such that $c_i^H(\GG)=(-1)^i\th^i/i!$, where $\th=c_1^H(\OO(\Theta))$. Then $(A,\Theta)$ is a Jacobian of
a smooth curve.
\end{thm}

Here is a version of this result that is better adapted to our context. 

\begin{prop}\label{Jac-prop} 
Let $(A,\Theta)$ be a principally polarized abelian variety, such that $\Theta$ is irreducible.
Assume that there exists a coherent sheaf $\FF$ on $A$, supported sheaf theoretically on $\Theta$ and of length $1$
at the general point of $\Theta$, 
such that
the Fourier transform of $\FF$ is supported in dimension $\le 1$. Then $(A,\Theta)$ is a Jacobian.
\end{prop}

\Pf . We will reduce this to Theorem \ref{MMG-thm}.
By assumption, $\FF=F(E)$, where $E$ is an object of derived category of $\hat{A}\simeq A$ with the
support $Z\sub A$ of dimension $\le 1$. Pick an ample divisor $D$ that intersects each component
of $Z$ transversally, so that all the maps $\und{H}^iE\to \und{H}^iE(D)$ are injective.
Consider the exact triangle
$$E\to E(D)\to E'\to E[1].$$
Applying the Fourier transform and passing to cohomology we get an exact sequence
$$0\to \und{H}^{-1}F(E(D))\to \und{H}^{-1}F(E')\to \FF\to \und{H}^0F(E(D))\to \und{H}^0F(E')\to 0$$
and isomorphisms
$$\und{H}^iF(E(D))\rTo{\sim} \und{H}^iF(E') \ \text{ for } i\neq 0,-1.$$
If $D$ is sufficiently ample then the sheaves $\und{H}^iE(D)$ and $\und{H}^iE'$ are IT-sheaves of index $0$. Hence,
we have 
$$\und{H}^iF(E(D))\simeq F(\und{H}^iE(D)), \ \ \und{H}^iF(E')\simeq F(\und{H}^iE').$$
Thus, $\und{H}^0F(E(D))$ is a vector bundle and the map $\FF\to H^0F(E(D))$ from the torsion sheaf $\FF$
is zero. Thus, from the above exact sequence we get a short exact sequence
$$0\to F(\und{H}^{-1}E(D))\to F(\und{H}^{-1}E')\to \FF\to 0.$$
On the other hand, since the map $\und{H}^0E\to \und{H}^0E(D)$ is injective, we obtain an exact sequence
$$0\to \und{H}^{-1}E\to \und{H}^{-1}E(D)\to \und{H}^{-1}E'\to 0,$$
so the sheaf $\und{H}^{-1}E'$ has zero-dimensional support.
This implies that $c_1(F(\und{H}^{-1}E'))=0$. Hence, the sheaf 
$$\GG:=F(\und{H}^{-1}E(D))$$ satisfies
$$c_1^H(\GG)=-c_1^H(\FF)=-\th.$$ 
Also, since $\GG$ is the Fourier transform of a sheaf supported in
dimension $1$, we have $\ch_{\ge 2}^H(\GG)=0$ (as a consequence of Mukai's formula \cite[Cor.\ 1.18]{Mukai-2}),
which is equivalent to $c_i^H(\GG)=c_1^H(\GG)^i/i!$.
Finally, since $\und{H}^{-1}E(D)$ is an IT-sheaf of index $0$, $\GG$ is a WIT-sheaf of index $g$. Thus, $\GG$ satisfies the
conditions of Theorem \ref{MMG-thm}. 
\ed


\begin{cor}
Conjecture B implies Conjecture A.
\end{cor}

\begin{proof}
We only need to check the ``if" part of Conjecture A. If there exists a coherent sheaf $\FF$ on $A$ with the stated properties
then by Conjecture B its Fourier transform has support of dimension $\le 1$. Hence, by Proposition \ref{Jac-prop}, $(A,\Th)$ is a Jacobian.
\end{proof}


Recall that the Jacobians of curves belong to the Andreotti-Mayer locus of principally polarized abelian varieties 
$(A,\Th)$ for which the singular locus of $\Th$ has dimension $\ge g-4$.

\begin{prop}\label{ch-2-prop} Assume $(A,\Theta)$ is a principally polarized abelian variety of dimension $g\ge 4$, with $\Theta$ irreducible.
Suppose there exists a coherent sheaf $\FF$ on $A$, sheaf theoretically supported on $\Theta$, of length $1$ at the general point of $\Theta$, 
and such that $\ch_2^H(\FF)=0$. 
Then $\dim\Sing(\Th)\ge g-4$.
\end{prop}

\Pf . Set $A':=A\setminus\Sing(\Th)$, $\Th':=\Th\setminus\Sing(\Th)$,
and let $N=\OO_{\Th}(\Th)$ be the normal line bundle to $\Th$.
Consider the closed embedding $i:\Th'\to A'$.
By assumption, $\FF|_{A'}=i_*F$, where $F$ is a coherent sheaf on $\Th'$, generically of rank $1$.
By the Grothendick-Riemann-Roch theorem applied to the embedding $i$, we have
$$\ch_2(\FF|_{A'})=i_*[\ch(F)\Td^{-1}(N|_{\Th'})]_2=i_*(c_1(F)-\frac{c_1(N|_{\Th'})}{2})$$
in $\CH_{g-2}(A')_{\Q}$. Assume that $\dim\Sing(\Th)<g-4$. Then the natural maps
\begin{equation}\label{Pic-restr-maps}
\Pic(A)\to \Pic(\Th)\to \CH_{g-2}(\Th)\to\CH_{g-2}(\Th')
\end{equation}
are isomorphisms (see \cite[Sec.\ 3.1]{BP}), so there exists a class $\ell\in\Pic(A)$ such that $c_1(F)$ is the
restriction of $\ell$ to $\Th'$. Thus, we get an equality in $\CH_{g-2}(A')_{\Q}$,
$$\ch_2(\FF|_{A'})=c_1(\Th)\cdot (\ell-\frac{c_1(\Th)}{2})|_{A'}.$$
Since the restriction map $\CH_{g-2}(A)\to\CH_{g-2}(A')$ is an isomorphism, this implies
that 
$$\ch_2(\FF)=c_1(\Th)\cdot (\ell-\frac{c_1(\Th)}{2})=0$$
in $\CH_{g-2}(A)_{\Q}$ and hence, we have the same equality in $H^4(A,\C)$. 
Since $\ch^H_2(\FF)=0$ in $H^4(A,\C)$ by assumption, we get
$$c_1^H(\Th)\cup (\cl(\ell)-\frac{c_1^H(\Th)}{2})=0,$$
where $\cl:\CH_{g-1}(A)\to H^2(A,\Z)$ is the cycle map. But the map
$$H^2(A,\C)\rTo{\cup c_1^H(\Th)} H^4(A,\C)$$
is injective being part of the Lefschetz $\ssl_2$-action (since $2<g$). Hence, 
$$\cl(\ell)-\frac{c_1^H(\Th)}{2}=0$$
in $H^2(A,\Z)$ which contradicts to the fact that $c_1^H(\Th)$ is a primitive vector in $H^2(A,\Z)$.
\ed

\begin{cor}\label{ch-2-cor}
Assume $(A,\Theta)$ is a principally polarized abelian variety of dimension $g\ge 4$, with $\Theta$ irreducible.
Suppose there exists a coherent sheaf $\FF$ on $A$, sheaf theoretically supported on $\Theta$, of length $1$ at the general point of $\Theta$, 
such that $\FF$, viewed as an object in $D^b(A)$ extends to a complex of vector bundles on the standard $1$-smooth thickening of $A$.
Then $\dim\Sing(\Th)\ge g-4$.
\end{cor}

\begin{proof}
Indeed, by Proposition \ref{ch-prop}, we have $\ch_{\ge 2}^H(\FF)=0$. It remains to apply Proposition \ref{ch-2-prop}.
\end{proof}

\section{Relation to deformation quantization}\label{quant-sec}

\subsection{Dg-constructions}\label{dg-constr-sec}

\subsubsection{NC-smooth thickenings}

Here we recall the construction of an NC-smooth thickening associated with a torsion free flat connection $\nabla$ on $X$
(see \cite{P-Tu}) and a parallel construction of deformation quantizations (due to Fedosov \cite{Fedosov}).

Namely, one can extend the connection $\nabla$ on $T_X$ to a derivation $D_1$ on 
$$\AA_X:=\Om^{\bullet}_X\ot_{\OO_X} \hat{T}_{\OO_X}(\Om^1_X),$$
extending the de Rham differential on $\Om^\bullet_X$ and such that for $\a\in \Om^1_X$ one has
$$D_1(1\ot \a)=\nabla(\a).$$
On the other hand, one has an $\Om^\bullet_X$-linear derivation $D_0$ of $\AA_X$ such that $D_0(1\ot \a)=\a\ot 1$.

In \cite{P-Tu} we studied the dg-algebra $(\AA_X,D)$, where $D=D_0+D_1$, and showed that $\und{H}^0(\AA_X,D)$ gives an NC-smooth thickening.
We also considered a more general version of this construction in \cite{P-Tu},
starting with a  twisted version of $\AA_X$ and a not necessarily flat connection $\nabla$, but we will not use this more general construction here.


\subsubsection{Deformation quantizations}

Recall that an (algebraic) deformation quantization of $X$ is a sheaf of associative
$k[\![h]\!]$-algebras $\DD$ on $X$, complete with respect to the $h$-adic topology and flat over $k[\![h]\!]$, equipped with
an isomorphism $\rho:\DD/h\DD\rTo{\sim}\OO_X$.
In particular, looking at the commutator in $\DD$ one gets a Poisson bracket on $\OO_X$ such that
$ab-ba=h\{\rho(a),\rho(b)\} \mod h^2\DD$, and one usually says that $\DD$ quantizes this Poisson
bracket. The Poisson bivector $P\in H^0(X,\bigwedge^2 T_X)$ is characterized by the condition
$\lan P, df\wedge dg\ran=\{f,g\}$, where $f$ and $g$ are local functions on $X$.

Now let us start with a Poisson bivector $P$ on $X$.
Assume that we have a torsion free flat connection $\nabla$ on $X$ such that the Poisson bivector 
$P\in H^0(X,\bigwedge^2 T_X)$ is $\nabla$-horizontal. This means
that for any vector field $v$ and any pair of functions $f,g$ we have
$$v\cdot \{f,g\}=\{v\cdot f,g\}+\{f, v\cdot g\}.$$

The sheaf of (possibly degenerate) formal Weyl algebras associated with $P$ is the quotient
$$\WW_P=\hat{T}(\Om^1)[\![h]\!]/\JJ_P,$$
where the two-sided ideal $\JJ_P$ is generated by
$$\om_1\ot\om_2-\om_2\ot\om_1- h\lan P, \om_1\wedge \om_2\ran.$$
By the PBW-theorem, we have a direct sum decomposition
\begin{equation}\label{Poisson-decomposition-eq}
\hat{T}(\Om^1)[\![h]\!]=\JJ_P\oplus \hat{S}(\Om^1)[\![h]\!]
\end{equation}
where $\hat{S}(\Om^1)$ consists of symmetric tensors.
Thus, we have a canonical isomorphism of $\OO_X[\![h]\!]$-modules
\begin{equation}\label{WP-Sym-eq}
\hat{S}(\Om_X^1)[\![h]\!]\rTo{\sim}\WW_P,
\end{equation}
whose reduction modulo $h$ is an isomorphism of algebras.

\begin{prop}\label{Moyal-prop}
(i) The decomposition of $\Om^\bullet\ot_\OO\hat{T}(\Om^1)[\![h]\!]$
induced by \eqref{Poisson-decomposition-eq}
is compatible with the the derivations $D_0$ and $D_1$ (extended $h$-linearly).

\noindent
(ii) Let $\ov{D}_0$ and $\ov{D}_1$ denote the induced derivations of $\Om^\bullet\ot\WW_P$.
Then $\DD_{P,\nabla}:=\und{H}^0(\Om^\bullet\ot\WW_P,\ov{D}_0+\ov{D}_1)$ is a deformation quantization of $P$.

\noindent
(iii) If $(x_1,\ldots,x_n)$ is a formal coordinate system at a point $p\in X$, such that $dx_i$ are $\nabla$-horizontal, then 
we have an identification of the completion of $\DD_{P,\nabla}$ at $p$ with $\C[\![x_1,\ldots,x_n,h]\!]$ with the Moyal product
\begin{equation}\label{Moyal-eq}
f*g=\sum_{k=0}^\infty \sum_{i_1,\ldots,i_k;j_1,\ldots,j_k}\frac{h^k}{k!2^k}\Pi_{i_1j_1}\ldots\Pi_{i_kj_k}(\pa_{i_1}\ldots\pa_{i_k}f)(\pa_{j_1}\ldots\pa_{j_k}g).
\end{equation}
\end{prop}

\begin{proof}
(i) We have $D_0(\om_1\ot\om_2-\om_2\ot\om_1)=0$, which easily implies that $D_0(\JJ_P)\sub \Om^1\ot \JJ_P$. On the other hand, for any $\a\in \Om^1$
one has $D_0(\a^{\ot n})\sub \Om^1\ot \hat{S}(\Om^1)$, hence $D_0(\hat{S}(\Om^1))\sub \Om^1\ot \hat{S}(\Om^1)$. 

To see invariance with respect to $D_1$,
let us choose a $\nabla$-horizontal basis $\a_1,\ldots,\a_n$ of $\Om_X^1$ in the formal neighborhood of some point. Then $\JJ_P$ is generated by
$\a_i\ot \a_j-\a_j\ot\a_i-h\Pi_{ij}$ where $\Pi_{ij}$ are constants. Since these elements are killed by $D_1$, we get $D_1(\JJ_P)\sub \Om^1\ot \JJ_P$.
Similarly, $S(\Om^1)$ is spanned by the monomials in $(\a_i)$, which are killed by $D_1$, so $D_1(\hat{S}(\Om^1))\sub \Om^1\ot \hat{S}(\Om^1)$.

\noindent
(ii),(iii) By (i), we have an isomorphism of complexes
$$(\Om^\bullet\ot \hat{S}(\Om^1)[\![h]\!],D_0^s+D_1^s)\rTo{\sim} (\Om^\bullet\ot \WW_P, \ov{D}_0+\ov{D}_1),$$
where $D_i^s$ is the restriction of $D_i$, $i=0,1$. Viewing $\hat{S}(\Om^1)$ as a completion of the symmetric algebra, it is easy to check that
$(\Om^\bullet\ot \hat{S}(\Om^1),D_0^s)$ is the completion of the Koszul complex ${\bigwedge}^\bullet (V)\ot S^\bullet(V)$ for $V=\Om^1$. Thus, we can
consider the standard homotopy 
$$H_i:{\bigwedge}^i(V)\ot S^j(V)\to {\bigwedge}^{i-1}(V)\ot S^{j+1}(V), \ i\ge 1,$$ 
of the positive degree terms of the Koszul complex given by the (suitably rescaled) action of the Euler vector field.
By \cite[Lem.\ 2.3.7]{P-Tu}, the operator $\id+HD_1^s$ extends to a well defined $\OO$-linear automorphism of $\Om^\bullet\ot \hat{S}(\Om^1)$ such that 
$$D_0^s+D_1^s=(\id+HD_1^s)^{-1}D_0^s(\id+HD_1^s),$$
which induces an isomorphism of $0$th cohomology we are interested in with the $0$th cohomology of $D_0^s$, which is identified $\OO_X[\![h]\!]$.

Explicitly, if $(x_1,\ldots,x_n)$ is a formal coordinate system as in (iii), then an element in $\ker(D_0^s+D_1^s)\sub\hat{S}(\Om^1)$ for a function $f$ is
$$\kappa(f):=f+\sum_{m\ge 1, i_1,\ldots,i_m} \frac{1}{m!}\pa_{i_1}\ldots\pa_{i_m}f\cdot dx_{i_1}\ot\ldots\ot dx_{i_m}.$$

It is well known (see e.g., \cite{Fedosov}) that if we equip $\hat{S}(\Om^1)$ using the identification \eqref{WP-Sym-eq} with $\WW_P$,
then $\kappa(f)\cdot \kappa(g)=\kappa(f*g)$, where $f*g$ is given by \eqref{Moyal-eq}. 
\end{proof}

We will refer to the deformation quantization $\DD_{P,\nabla}$ above as the {\it Moyal deformation quantization} associated with $(P,\nabla)$.

\subsection{Homomorphisms from NC-smooth thickenings to deformation quantizations}

It is natural to ask whether there exist any homomorphisms from NC-smooth thickenings of $X$ to deformation quantizations of $X$.
The answer is always positive in the affine case as the following result shows. Recall that in the affine case an NC-smooth thickening always exists and
is unique up to isomorphism 
(see \cite{Kapranov}).

\begin{prop}\label{NC-to-quant-affine-prop} 
Let $X=\Spec(A)$ be a smooth affine variety, and let $A^{NC}$ be the NC-smooth completion of $A$.
Then for any deformation quantization $D$ of $A$, there exists a homomorphism $A^{NC}\to D$ compatible with the filtrations
$(F^nA^{NC})$ and $(h^nD)$, and inducing 
the identity $A^{NC}/F^1=A\to A=D/hD$. For any such homomorphism, the induced map
$$\Om^2_A\simeq F^1/F^2\to hD/h^2D=hA$$
is given by the Poisson bivector associated with $D$.
\end{prop}

\begin{proof}
It suffices to construct recursively a collection of compatible homomorphisms $\phi_n:A^{NC}/F^nA^{NC}\to D/h^nD$, starting with the identity map
$A^{NC}/F^1\to D/hD$. Assume we already constructed a homomorphism $\phi_{n-1}:A^{NC}/F^{n-1}A^{NC}\to D/h^{n-1}D$ compatible with the filtrations.
Now we use the fact that $A^{NC}/F^nA^{NC}$ is the universal square zero extension of $A^{NC}/F^{n-1}A^{NC}$ by a central bimodule (see
\cite[Prop.\ (1.6.2)]{Kapranov}). This implies that $A^{NC}/F^nA^{NC}$ maps to any other central extension of $A^{NC}/F^{n-1}A^{NC}$ (see
\cite[Prop.\ (1.3.8)]{Kapranov}).
Now $D/h^nD$ is a central extension of $D/h^{n-1}D$ by $D/hD$, and we can pull it back to get a central extension of $A^{NC}/F^{n-1}A^{NC}$. 
Applying the universality we get the required homomorphism $\phi_n$.

The last assertion is obtained by considering the commutator maps 
$$F^0/F^1\times F^0/F^1\to F^1/F^2, \ \ D/hD\times D/hD\to hD/h^2D,$$
and observing that the first one is identified with $(f,g)\mapsto df\we dg$ and the second is the Poisson bracket associated with $D$.
\end{proof}

Now we consider the global setting for the NC-thickening associated with a torsion-free flat connection.

\begin{prop}\label{NC-to-quant-prop} 
Let $\OO_X^{NC}$ be the smooth NC-thickening of $X$ associated with a torsion-free flat connection 
$\nabla$ on $X$. 
Let $P$ be a $\nabla$-horizontal Poisson bivector on $X$,
and let $\DD=\DD_{P,\nabla}$ be the Moyal deformation quantization algebra associated with $(P,\nabla)$ (see Proposition \ref{Moyal-prop}).
Then there exists a 
homomorphism
\begin{equation}\label{NC-to-def-eq}
\OO_X^{NC}\to \DD,
\end{equation}
compatible with the filtrations $(F^n=F^n\OO_X^{NC})$ and $(h^n\DD)$, such that the induced map $\OO_X=F^0/F^1\to \DD/h\DD=\OO_X$ is the identity,
and the induced map $\Om^2_X=F^1/F^2\to h\DD/h^2\DD=h\cdot \OO_X$ is given by $h\Pi$.
\end{prop}

\begin{proof} We use the dg-constructions of Sec.\ \ref{dg-constr-sec} defining $\OO_X^{NC}$ and $\DD_{P,\nabla}$.
The natural projection gives a homomorphism of sheaves of dg-algebras
$$(\Om_X^\bullet\ot_{\OO_X}\hat{T}(\Om^1_X), D_0+D_1)\to (\Om_X^\bullet\ot_{\OO_X}\WW_P, \ov{D}_0+\ov{D}_1).$$
Passing to the $0$th cohomology sheaves we get homomorphism \eqref{NC-to-def-eq}.
The identification of $F^0/F^1$ (resp., $\DD/h\DD$) with $\OO_X$ is induced by the natural projection $\hat{T}(\Om^1)\to \OO_X$ 
(resp., $\WW_P\to \WW_P/h\WW_P\to \OO_X$). This implies the assertion about the induced map on $F^0/F^1$. The assertion about the induced map
on $F^1/F^2$ follows by considering commutators, as in the proof of Proposition \ref{NC-to-quant-affine-prop}.
\end{proof}


\begin{cor} Let $E$ is an object of $D^b(A)$, where $A$ is an abelian variety, such that the Fourier-Mukai transform of $E$ has support of dimension $\le 1$.
Then $E$ extends to a perfect complex over the Moyal deformation quantization of any constant Poisson bracket $P$ on $A$.
\end{cor}

\begin{proof}
By Theorem \ref{NC-FM-thm}, there exists an extension of $E\in D^b(A)$ to a perfect complex $\wt{E}$ of $\OO_A^{NC}$-modules.
Now consider a homomorphism from $\OO_A^{NC}$ to the Moyal quantization $\AA_P$ of $\OO_A$ associated
with $P$. Then $E_P=\AA_P\ot_{\OO_A^{NC}} E$ is a perfect complex of $\AA_P$-modules extending $E$.
\end{proof}

\begin{exs} 1. Let $C$ be a smooth genus $2$ curve, and let $i:C\to J$ denote a standard embedding into the Jacobian. Since $J$ is $2$-dimensional, there is a unique nonzero constant Poisson bracket $P$ on $J$, up to rescaling. Then for a line bundle $L$ on $C$, the coherent sheaf $i_*L$ extends to a module over the Moyal deformation quantization $\AA_P$ of $P$,
flat over $k[\![h]\!]$, if and only if
$\deg(L)=1$ (by the main result of \cite{BGKP}). It is easy to see that this is equivalent to the condition that the Fourier transform of $i_*L$ has support of dimension $\le 1$.
Thus, by Theorem \ref{NC-FM-thm}, we get that $L$ extends to a flat module over $\AA_P$ if and only if it extends to a perfect complex over $\OO_J^{NC}$.

\noindent
2. More generally, let $A$ be an abelian surface, and let $i:C\hra A$ be a smooth connected divisor, which is a curve of genus $g\ge 2$, and let $L$ be a line bundle on $C$.
Then still by \cite{BGKP}, the sheaf $i_*L$ extends to the Moyal deformation quantization of a nonzero constant Poisson bracket $P$ on $A$,
flat over $k[\![h]\!]$, if and only if $\deg(L)=g-1$.
Assuming that $\deg(L)=g-1$, we get that the Fourier transform of $L$ on the Jacobian $J_C$ has form $\FF[-1]$, where $\FF$ is a coherent sheaf supported sheaf-theoretically
on the theta-divisor $\Th_L\sub J_C$ and of generic rank $1$ on $\Th_L$. The Fourier transform of $i_*(L)$ on $\hat{A}$ is $Lf^*\FF[-1]$, where $f:\hat{A}\to J_C$
is induced by $i$. Thus, if $f(\hat{A})$ is not contained in $\Th_L$, this Fourier transform is supported in dimension $\le 1$, which implies by Theorem \ref{NC-FM-thm} that $i_*L$ extends to the standard NC-smooth thickening of $A$. 
I do not know examples when $f(\hat{A})$ is contained in $\Th_L$ (see however \cite{AC} and references therein): if such examples exist then checking whether $i_*L$ extends to $A^{NC}$ would be a good testing case for Conjecture B. 
\end{exs}

\subsection{First order deformation quantizations of abelian varieties and the theta divisor}

One can also ask whether a  version of the Conjecture A holds in which the NC-thickening
of $A$ is replaced by the deformation quantization with respect to a generic constant Poisson structure on $A$.
We are going to show that the condition of extendability to
the 1st order deformation quantization implies that $(A,\Th)$ is in the Andreotti-Mayer locus.

Let $P\in H^0(X,\bigwedge^2 T_X)$ be a global bivector field on a smooth variety $X$. 
By the {\it standard 1st order deformation of $X$ associated with $P$} we mean the sheaf of algebras
$\OO^P_{\le 1}=\OO_X[h]/h^2$ with the multiplication 
$$(f_0+f_1h)\star (g_0+g_1h)=f_0g_0+(\{f_0,g_0\}+f_0g_1+f_1g_0)h,$$
where $\{f_0,g_0\}=\lan P, df_0\wedge dg_0\ran$.

\begin{prop}\label{1st-order-def-qu-prop} 
Let $(A,\Theta)$ be a principally polarized abelian variety of dimension $g\ge 4$, such that $\Theta$ is irreducible. There exists a countable union $Z$ of proper linear subspaces in 
$H^0(A,\bigwedge^2 T_A)$, such that any bivector $P\in H^0(A,\bigwedge^2 T_A)\setminus Z$
has the following property. 
Suppose for some line bundle $L$ on $\Th':=\Th\setminus \Sing(\Th)$ the coherent sheaf $i_*L$ on $A':=A\setminus\Sing(\Th)$, where $i:\Th'\to A'$ is the embedding,
extends to a module over the standard 1st order deformation of $A'$ associated with $P$, flat over $k[h]/h^2$. Then $\dim\Sing(\Th)\ge g-4$.
\end{prop}

\Pf . Assume $\dim\Sing(\Th)<g-4$ and let us show that this leads to a contradiction.
Since any divisor is coisotropic, the morphism
$\eta=\eta_P:\Om^1_A\to T_A$ associated with $P$ induces a morphism
$\eta_{\Th}:\Om^1_{\Th}\to N:=\OO_\Th(\Th)$, so that we have a commutative  diagram
\begin{equation}\label{forms-comm-diag}
\begin{diagram}
\Om^1_A &\rTo{\eta}& T_A\\
\dTo{} &&\dTo{}\\
\Om^1_\Th &\rTo{\eta_\Th}& N
\end{diagram}
\end{equation}
Let us also set $N':=N|_{\Th'}$, $\eta_{\Th'}:=\eta_{\Th}|_{\Th'}$.

We will use the result of Baranovsky-Ginzburg-Pecharich \cite[Thm.\ 7]{BGP} which states
that the existence of the extension in question implies the following equation in $H^1(\Th',N')$:
\begin{equation}\label{BGP-constraint}
\eta_{\Th'} [-c_1^H(N')+2c_1^H(L)]=0,
\end{equation}
where we view first Chern classes as elements in $H^1(\Th',\Om^1_{\Th'})$.
Since $\dim\Sing(\Th)<g-4$, the restriction maps \eqref{Pic-restr-maps} are isomorphisms, so the line bundle $L$ is a restriction of some line bundle $\wt{L}$ on $A$.
Thus, 
$$-c_1^H(N')+2c_1^H(L)=i^*(-c_1^H(\OO_A(\Th)|_{A'})+2c_1^H(\wt{L}|_{A'})),$$
where 
$$i^*:H^1({A'},\Om^1_{A'})\to H^1(\Th',\Om^1_{\Th'})$$
is the natural pull-back map. Using the restriction of the diagram \eqref{forms-comm-diag} to ${A'}$ we 
see that the composition of this pull-back map with the map
$$H^1(\Th',\Om^1_{\Th'})\rTo{\eta_{\Th'}} H^1(\Th',N')$$
is equal to the composition
$$H^1({A'},\Om^1_{A'})\rTo{\eta} H^1({A'},T_{A'})\to H^1(\Th',N').$$
Hence, the equation \eqref{BGP-constraint} means that the composition
\begin{equation}\label{long-composition-eq}
H^1(A,\Om^1_A)\to H^1({A'},\Om^1_{A'})\rTo{\eta|_{A'}} H^1({A'},T_{A'})\to H^1(\Th',N')
\end{equation}
maps the class $-c_1^H(\OO(\Th))+2c_1^H(\wt{L})$ to zero.
Since the class $c_1^H(\Th)\in H^2(A,\Z)$ is primitive, if we prove that for $\eta=\eta_P$ associated with a sufficiently generic  $P$, 
the composition \eqref{long-composition-eq} restricted to the subgroup of Chern classes of line bundles is injective, this will give a contradiction.
 
We can rewrite \eqref{long-composition-eq} as
$$H^1(A,\Om^1_A)\rTo{\eta} H^1(A,T_A)\to H^1(\Th, N)\to H^1(\Th',N').$$
Next we observe that since $\Th$ is Cohen-Macaulay and $\Sing(\Th)$ has codimension $\ge 3$ in $\Th$, 
we have the vanishing of the local cohomology, $\und{H}_{\Sing(\Th)}^i(N)=0$ for $i\le 2$. This
implies that the map $H^1(\Th,N)\to H^1(\Th',N^{reg})$ is an isomorphism.
Thus, it is enough to prove injectivity of the composition
$$\NS(A)\to H^1(A,\Om^1_A)\rTo{\eta} H^1(A,T_A)\to H^1(\Th,N)$$
for sufficiently generic $P$. From the exact sequence
$$0\to \OO_A\to \OO_A(\Th)\to N\to 0$$
we get a connecting homomorphism $H^1(\Th,N)\to H^2(A,\OO_A)$. It is enough to prove injectivity after post-composing with this map.

By Lemma \ref{divisor-class-lem} below, the composition 
$$H^1(A,T_A)\to H^1(\Th,N)\to H^2(A,\OO_A)$$
is given by the product with $c_1^H(\Th)\in H^1(A,\Om^1_A)$. Therefore, we are reduced to checking
injectivity of the composed map
\begin{equation}\label{NS-to-H2-map}
\NS(A)\to H^1(A,\Om^1_A)\rTo{\eta} H^1(A,T_A)\rTo{\cup c_1^H(\Th)} H^2(A,\OO)
\end{equation}
for sufficiently generic $P$. 
Note that the last arrow can be written as the composition
$$H^1(A,T_A)=T_{A,0}\ot H^1(A,\OO)\rTo{\phi_\Th\ot \id} 
T_{\hat{A},0}\ot H^1(A,\OO)=H^1(A,\OO)^{\ot 2}\to 
{\bigwedge}^2 H^1(A,\OO)=H^2(A,\OO),$$
where $\phi_\Th: T_{A,0}\to T_{\hat{A},0}$ is the tangent map to the principal polarization isomorphism
$A\to\hat{A}$ associated with $\Th$. Thus, a class $\la\in\NS(A)$ is in the kernel of \eqref{NS-to-H2-map} 
if and only if 
\begin{equation}\label{eta-lambda-in}
\eta(\la)\in (\phi_\Th\ot\id)^{-1}(S^2H^1(A,\OO))\sub T_{A,0}\ot H^1(A,\OO).
\end{equation}
Since $\NS(A)$ is a finitely generated abelian group, it remains to check that for every nonzero 
$\la\in\NS(A)$ the condition \eqref{eta-lambda-in} defines a proper linear subspace of $\eta$'s, where we think of
$\eta$ as a skew-symmetric map $T^*_{A,0}\to T_{A,0}$.

Set $V:=T_{A,0}$.
Recall that there is a natural identification $H^1(A,\OO)=\bar{V}^*$ (the dual of the complex conjugate
vector space to $V$), so that with every $\la\in\NS(A)$ the corresponding element
$$c_1(\la)\in H^1(A,\Om^1_A)=V^*\ot H^1(A,\OO)=V^*\ot\bar{V}^*$$
is represented by a Hermitian form $H=H_\la$ associated with $\la$, that is $H(v_1,v_2)$ is $\C$-linear in 
$v_1$, $\C$-antilinear in $v_2$ and satisfies $H(v_2,v_1)=H(v_1,v_2)$.
Let $H_0(v_1,v_2)$ be the Hermitian form associated with the class of $\Theta$. Then we can write
any other Hermitian form uniquely as $H(v_1,v_2)=H_0(Qv_1,v_2)$, for a $\C$-linear operator $Q:V\to V$ which is 
self-adjoint with respect to $H_0$. Also, we can use the identification $\bar{V}\to V^*$ given by $H_0$ to 
view the operator $\eta:V^*\to V$ as a ($\C$-linear) operator $\psi:\bar{V}\to V$. Then the
condition \eqref{eta-lambda-in} is equivalent to the condition that the form $H(\psi v_1,v_2)$ is symmetric
(where $H=H_\la$). In terms of the operator $Q$ this condition becomes
$Q\psi=-\psi\ov{Q}$. If we choose a $\C$-basis of $V$, orthonormal with respect to $H_0$ then the matrices
corresponding to $Q$ and $\psi$ satisfy $\ov{Q}=Q^T$ and $\psi^T=-\psi$. Thus, the condition
\eqref{eta-lambda-in} is equivalent to $Q\psi=-\psi Q^T$. It remains to
observe that for every nonzero matrix $n\times n$-matrix $Q$ (where $n>2$), there exists a skew-symmetric matrix $\psi$ 
such that $Q\psi+\psi Q^T\neq 0$ (this can be checked by taking $\psi$ of the form $E_{ij}-E_{ji}$
where $E_{ij}$ are elementary matrices).
\ed

\begin{lem}\label{divisor-class-lem} 
Let $D\sub X$ be an effective divisor in a smooth variety. Then the composition 
$$H^1(X,\Om^1_X)\to H^1(X,\OO_D(D))\to H^2(X,\OO),$$
where the second map comes from the exact sequence $0\to \OO_X\to \OO_X(D)\to \OO_D(D)\to 0$,
is equal to the product with the class $c_1^H(D)\in H^1(X,\Om^1_X)$.
\end{lem}

\Pf . We have a commutative diagram
\begin{diagram}
0\to &\OO_X &\rTo{}& D_{\le 1}(\OO_X(D))&\rTo{}& T_X&\to 0\\
&\dTo{\id}&&\dTo{}&&\dTo{}\\
0\to &\OO_X &\rTo{}& \OO(D)&\rTo{}& \OO_D(D)&\to 0\\
\end{diagram}
where $D_{\le 1}(\OO_X(D))$ is the sheaf of differential operators $\OO_X(D)\to \OO_X(D)$ of order $\le 1$,
and the middle vertical arrow sends an operator $D$ to $D\cdot 1$, where $1\in\OO_X\sub\OO_X(D)$.
It remains to use the fact that the class of the extension given by the first row of the diagram is
exactly $c_1^H(D)$.
\ed

\begin{rem} There is a homomorphism $\OO^{NC}_{\le 1}\to\OO^P_{\le 1}$
from the standard 1-smooth thickening 
to the 1st order deformation associated with any bivector $P$. If we start with a module of finite projective dimension over $\OO^{NC}_{\le 1}$ by extending scalars
we get a module $M$ of finite projective dimension
over $\OO^P_{\le 1}$. We claim that any such module is automatically flat
over $k[h]/h^2$. Indeed, we have to check that the infinite periodic complex 
$$\ldots\to M\rTo{h} M\rTo{h}M\to\ldots$$
is exact, and this follows immediately from the fact that $M$ has a finite resolution by modules for which this 
exactness holds.
Thus, from Proposition \ref{1st-order-def-qu-prop} we get another proof of Corollary \ref{ch-2-cor} for $\FF$ of the form $j^{ns}_*L$ (as in
Theorem \ref{Jac-thm}).
\end{rem}




\subsection{Example with the standard $1$-smooth thickening of an abelian surface}
\label{surface-ex-sec}

For every abelian surface $A$ we will give an example of an object $E\in D^b(A)$ such that
the Fourier transform of $E$ has full support but $E$ extends to the standard 1-smooth thickening of $A$.
This shows that in Conjecture B it is not to sufficient to look at the first order NC-thickening.

First, we observe that trivializing $\om_A$, we can view the standard $\om_A$-valued bracket $\{f,g\}=df\we dg$ as a Poisson bracket $\Pi$ on $\OO_A$.
Hence, we have the identification of the sheaf $\OO_A^{1-NC}$
with the $1$st order deformation of $\OO_A$ associated with $\Pi$.

\begin{lem}\label{H2-lem} 
Let $\a\in H^2(A,\OO_A)$ be a nonzero element. 
Let $F:=\Cone(\a:\OO_A\to \OO_A[2])$. Then the morphism $\a:F\to F[2]$ is zero in $D^b(A)$.
\end{lem}

\begin{proof}
Working with a minimal $A_\infty$-category corresponding to $D^b(A)$, we can represent $F$ by the twisted complex $[\OO_A\rTo{\a}\OO_A[1]]$
(where the differential has degree $1$). Then the morphism $\a$ is represented by an obvious chain map 
$$[\OO_A\rTo{\a}\OO_A[1]]\to [\OO_A[2]\rTo{\a}\OO_A[3]]$$
with the components given by $\a$ (there can be no nontrivial component $\OO_A\to \OO_A[3]$). But this chain map is homotopic to zero,
which implies the assertion.
\end{proof}


Our proof will use the extension of Fourier-Mukai transform to the equivalence between derived categories of the first order deformations of $D^b(A)$
constructed in \cite{Toda} (it has been extended to certain higher order deformations in \cite{BBP}).

\begin{prop} Let $e\in A$ be the neutral element, and let $u\in \Ext^2_A(\OO_e,\OO_e)$ denote a nonzero element.
Then the object $G:=\Cone(u:\OO_e\to \OO_e[2])$ of $D^b(A)$ extends to a perfect complex in $D^b(A^{1-NC})$.
\end{prop}

\begin{proof}
By the result of Toda, \cite[Thm.\ 1.1]{Toda}, it is enough to prove that the Fourier transform $F$ of $G$ extends to a perfect object of the derived category $D^b(\Coh^\a(A))$,
where $\Coh^\a(A)$ is the first order deformation of $\Coh(A)$ associated with a nonzero class $\a\in H^2(A,\OO_A)$. Furthermore, by \cite[Prop.\ 6.1]{Toda}, it is enough
to check that the morphism $\a:F\to F[2]$ vanishes. But this is proved in Lemma \ref{H2-lem}.
\end{proof}

\end{document}